# Finitary conditions for graph products of monoids

### v3

@@ 3. The ascending chain condition on principal left ideals
$z]\in\mathscr{GP}[a_{1}\circ\cdots\circ a_{n}]=\mathscr{GP}[a].$
@@
$\mathscr{GP}[a_{1}\circ\cdots\circ a_{n}]=\bigcap_{1\leq i\leq n}\mathscr{GP}[a_{i}]=\bigcap_{1\leq i\leq n}\mathscr{GP}[b_{i}]=\mathscr{GP}[b_{1}\circ\cdots\circ b_{n}]=\mathscr{GP}[b].$
@@
is in left Foata normal form with block length $\lambda(i)$ and blocks $p^{i}_{j}\in X^{*}$, $1\leq j\leq\lambda(i)$, and that the first block contains $p^{i}_{1}$ contains no left invertible letters.
@@ 4. Weakly left noetherian
For the converse, suppose that each $M_{i}$ is weakly left noetherian, and $J:=\{i\in I:M_{i}\mbox{ is not a group}\}$ is finite. Then $P\cong P_{J}\times P_{G}$, where $P_{J}$ is the direct product of the monoids in $\mathcal{M}_{J}:=\{M_{j}:j\in J\}$ and $P_{G}$ is the restricted direct product of the groups in $\mathcal{M}_{G}:=\{M_{j}:j\in I\setminus J\}$. Certainly $P_{2}$ is a group, and so is weakly left noetherian. The result then follows from Lemma 4.2. ∎
@@ 5. Reduction of products
where $a^{K+1}$ is $a^{K}$ with $a_{o}$ deleted. We extend our injection $\theta$ to $[1,K+1]$ by defining $o=(K+1)\theta$. If $ya_{o}\notin I$ then we set $i_{h+1}=K+1$ and put $y=s_{i_{h+1}\theta}$ so that $ya_{o}=s_{i_{h+1}\theta}a_{(K+1)\theta}\notin I$; this is ‘glueing move’. If $ya_{o}\in I$ then we set $j_{k+1}=K+1$ and put $y=s_{j_{k+1}\theta}$ so that $ya_{o}=s_{j_{k+1}\theta}a_{(K+1)\theta}\in I$; this is ‘deletion move’. In the case of a glueing move we have increased the cardinality of $I$ by one to obtain $I^{\prime}$; then $h+1,k,I^{\prime},J,\theta,s^{K+1}$ and $a^{K+1}$ are such that conditions (1)-(5) hold. In the case of a deletion move we have increased the cardinality of $J$ by one to obtain $J^{\prime}$; then $h,k+1,I,J^{\prime},\theta,s^{K+1}$ and $a^{K+1}$ are such that conditions (1)-(5) hold.
@@
$a^{\prime}\circ b\equiv a^{(\ell)}\circ a^{(\ell)}\circ w_{a,b}\circ w\equiv b^{\prime}\circ a$
@@ 6. Left ideal Howson
    $a$ shuffles to $a_{1}\circ\cdots\circ a_{M}\circ a^{\prime}$;
@@
$s^{\prime}]=[w\circ d^{\prime}],[t^{\prime}]=[w\circ c^{\prime}]\mbox{ and }[d^{\prime}\circ c]=[c^{\prime}\circ d].$
@@
We return to the reduced and equivalent words $d^{\prime}\circ c$ and $c^{\prime}\circ d$. We will employ Lemma 5.9 with $c$ and $d$ taking the role of $a$ and $b$. To that end, let
$c^{\ell}=s_{1}a_{1}\circ\cdots\circ s_{K},\,c^{r}=a^{M},\,d^{\ell}=t_{1}b_{1}\circ\cdots\circ t_{N}\mbox{ and }d^{r}=b^{N}.$
We let $I_{c},J_{c},I_{d}$ and $J_{d}$ be defined as in that lemma. Then
@@
Let $a^{(\ell)}$ be the reduced word with the same support as $c^{(\ell)}$, but replacing each $s_{i}a_{i}$ by $a_{i}$; let $b^{(\ell)}$ be defined similarly, and let $w^{\prime}=a^{(\ell)}\circ b^{(\ell)}\circ w$. Note that $w^{\prime}$ consists only of letters of $a$ and $b$. Let $\sigma:I_{a}\rightarrow I_{b}$ be defined as in Lemma 5.9 so that $s_{o}a_{o}=t_{o\sigma}b_{o\sigma}$. We now call upon the fact that all the vertex monoids are left ideal Howson, and we let $X_{o}$ be a finite set of generators of the left ideal $M_{o}a_{o}\cap M_{o}b_{o\sigma}$ for all $o\in I_{a}:=\{o_{1},\cdots,o_{s}\}$, where $o=\mbox{$\mathbf{s}$}(a_{o})$. We claim that $X$ is a finite set of generators for $\mathscr{I}$, where
$X=\{[x_{o_{1}}\circ\cdots\circ x_{o_{s}}\circ w^{\prime}]:x_{o}\in X_{o}\}.$
@@
First, we show that $\mathscr{GP}X\subseteq\mathscr{I}$. Let $x=x_{o_{1}}\circ\cdots\circ x_{o_{s}}\circ w^{\prime}\in X$. For $1\leq i\leq s$ choose $p_{o_{i}},q_{o_{i}\sigma}\in M_{o_{i}}$ such that $x_{o_{i}}=p_{o_{i}}a_{o_{i}}=q_{o_{i}\sigma}b_{o_{i}\sigma}$. We let $\overline{d}=b^{(\ell)}\circ d^{(r)}$ and $\overline{c}=a^{(\ell)}\circ c^{(r)}$ so that $\overline{d}$ and $\overline{c}$ have the same support skeleton as $d$ and $c$ respectively.
@@
We form a word $s_{1}^{\prime}\circ\cdots\circ s_{M}^{\prime}$ with the same support skeleton as $s_{1}\circ\cdots\circ s_{M}$ as follows. If $o\in I_{a}$ then put $s_{o}^{\prime}=p_{o}$; if $o\in J_{a}$ we put $s_{o}^{\prime}=I_{o}$, where $s_{o}\in M_{o}$; if $o\in K_{a}$ let $s_{o}^{\prime}=s_{o}$. Similarly, we form a word $t_{1}^{\prime}\circ\cdots\circ t_{N}^{\prime}$ with the same support skeleton as $t_{1}\circ\cdots\circ t_{N}$ as follows. If $o\sigma\in I_{b}$ then put $t_{o\sigma}^{\prime}=q_{o\sigma}$; if $o\in J_{b}$ we put $t_{o}^{\prime}=I_{o}$, where $t_{o}\in M_{o}$; if $o\in K_{b}$ let $t_{o}^{\prime}=t_{o}$. Then
$\bar{d}\circ s_{1}^{\prime}\circ\cdots\circ s_{M}^{\prime}\circ a]=[\bar{c}\circ t_{1}^{\prime}\circ\cdots\circ t_{N}^{\prime}\circ b]\in\mathscr{I}.$
Moreover, with $p=x_{o_{1}}\circ\cdots\circ x_{o_{s}}$ we have
$x]=[x_{o_{1}}\circ\cdots\circ x_{o_{s}}\circ w^{\prime}]=[p][a_{o_{1}}\circ\cdots\circ a_{o_{s}}\circ a^{(\ell)}\circ b^{(\ell)}\circ w]=[\bar{d}\circ s_{1}^{\prime}\circ\cdots\circ s_{M}^{\prime}\circ a]$
so that $x]\in\mathscr{I}$ and consequently, $\mathscr{GP}X=\mathscr{I}$.
@@
Now we show that $\mathscr{GP}X\supseteq\mathscr{I}$. To this end, let us return to an arbitrary $z]\in\mathscr{I}$. Then $z]=[s\circ a]=[t\circ b]$ where $s$ and $t$ are obtained as at the start of the proof. Note that it was the existence of any such $[z]$ that determined the support skeleton of $d^{\prime}\circ c$ and $c^{\prime}\circ d$; the support skeleton does not depend up the choice of $z$ and by Remark 5.8 is unique. For $o_{i}\in I_{a}$ we have that $s_{o_{i}}a_{o_{i}}=t_{o_{i}\sigma}b_{o_{i}\sigma}$ so that $s_{o_{i}}a_{o_{i}}=k_{o_{i}}x_{o_{i}}$ for some $x_{o_{i}}\in X_{o_{i}}$. Put $x_{o_{i}}=p_{o_{i}}a_{o_{i}}$ as above for $1\leq i\leq s$ and let $x=x_{o_{1}}\circ\cdots\circ x_{o_{s}}\circ w^{\prime}\in X$. Now list the elements of $J_{a}$ and $J_{b}$ respectively as
$J_{a}=\{e_{1},\cdots,e_{t}\}\mbox{ and }J_{b}=\{f_{1},\cdots,f_{r}\}.$
@@
$s\circ a]=[w\circ(k_{o_{1}}\circ\cdots\circ k_{o_{s}}\circ s_{e_{1}}\circ\cdots\circ s_{e_{t}}\circ t_{f_{1}}\circ\cdots\circ t_{f_{r}})][x]\in\mathscr{GP}X.$
@@ 4. $\bullet$
We call the pair of reductions $(\theta,\psi)$ a parallel pair and say that $([s],[t])\in\mathbf{l}([a])$ belong to a $(\theta,\psi)$ reduction. Note that there are only finitely many types of parallel pairs in $\mathbf{l}([a])$, since there are only finitely many reduction functions for $a$.
@@
For a parallel pair of reduction functions $(\theta,\psi)$, and $B^{\prime},A^{\prime}$ defined as above we let $B^{\prime}_{r}$ be $B^{\prime}$ where we replace any $t_{u_{k}\psi}a_{u_{k}\psi}$ in $B_{1}$ by $a_{u_{k}\psi}$ (so that $B_{r}^{\prime}$ is a product of letters of $a$) and let $A^{\prime}_{r}$ be $A^{\prime}$ where we replace any $s_{i_{l}\theta}a_{i_{l}\theta}$ in $B_{1}$ by $a_{i_{l}\theta}$ (so that $A_{r}^{\prime}$ is a product of letters of $a$). We then set
@@
We have that $(z^{\theta,\psi}_{\theta}],[z^{\theta,\psi}_{\psi}])\in\mathbf{l}([a])$.

### v4

@@ 3. The ascending chain condition on principal left ideals
$z]\in\mathscr{GP}[a_{1}\circ\cdots\circ a_{n}]=\mathscr{GP}[a].\qed$
@@
$\mathscr{GP}[a_{1}\circ\cdots\circ a_{n}]=\bigcap_{1\leq i\leq n}\mathscr{GP}[a_{i}]\subseteq\bigcap_{1\leq i\leq n}\mathscr{GP}[b_{i}]=\mathscr{GP}[b_{1}\circ\cdots\circ b_{n}]=\mathscr{GP}[b].$
@@
is in left Foata normal form with block length $\lambda(i)$ and blocks $p^{i}_{j}\in X^{*}$, $1\leq j\leq\lambda(i)$, and that the first block $p^{i}_{1}$ contains no left invertible letters.
@@ 4. Weakly left noetherian
For the converse, suppose that each $M_{i}$ is weakly left noetherian, and $J:=\{i\in I:M_{i}\mbox{ is not a group}\}$ is finite. Then $P\cong P_{J}\times P_{G}$, where $P_{J}$ is the direct product of the monoids in $\mathcal{M}_{J}:=\{M_{j}:j\in J\}$ and $P_{G}$ is the restricted direct product of the groups in $\mathcal{M}_{G}:=\{M_{j}:j\in I\setminus J\}$. Certainly $P_{G}$ is a group, and so is weakly left noetherian. The result then follows from Lemma 4.2. ∎
@@ 5. Reduction of products
where $a^{K+1}$ is $a^{K}$ with $a_{o}$ deleted. We extend our injection $\theta$ to $[1,K+1]$ by defining $o=(K+1)\theta$. If $ya_{o}\notin I$ then we set $i_{h+1}=K+1$ and put $y=s_{i_{h+1}\theta}$ so that $ya_{o}=s_{i_{h+1}\theta}a_{(K+1)\theta}\notin I$; this is a ‘glueing move’. If $ya_{o}\in I$ then we set $j_{k+1}=K+1$ and put $y=s_{j_{k+1}\theta}$ so that $ya_{o}=s_{j_{k+1}\theta}a_{(K+1)\theta}\in I$; this is a ‘deletion move’. In the case of a glueing move we have increased the cardinality of $I$ by one to obtain $I^{\prime}$; then $h+1,k,I^{\prime},J,\theta,s^{K+1}$ and $a^{K+1}$ are such that conditions (1)-(5) hold. In the case of a deletion move we have increased the cardinality of $J$ by one to obtain $J^{\prime}$; then $h,k+1,I,J^{\prime},\theta,s^{K+1}$ and $a^{K+1}$ are such that conditions (1)-(5) hold.
@@
$a^{\prime}\circ b\equiv a^{(\ell)}\circ b^{(\ell)}\circ w_{a,b}\circ w\equiv b^{\prime}\circ a$
@@ 6. Left ideal Howson
    $a$ shuffles to $a_{1}\circ\cdots\circ a_{M}\circ a^{M}$;
@@
$s^{\prime}]=[u\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\circ d^{\prime}],[t^{\prime}]=[u\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\circ c^{\prime}]\mbox{ and }[d^{\prime}\circ c]=[c^{\prime}\circ d]$
for some $u\in X^{*}$.
@@
We return to the reduced and equivalent words $d^{\prime}\circ c$ and $c^{\prime}\circ d$. We will employ Lemmas 5.7 and 5.9 with $c$ and $d$ taking the role of $a$ and $b$. To that end, let
$c^{\ell}=s_{1}a_{1}\circ\cdots\circ s_{M}a_{M},\,c^{r}=a^{M},\,d^{\ell}=t_{1}b_{1}\circ\cdots\circ t_{N}b_{N}\mbox{ and }d^{r}=b^{N}.$
We let $j_{1},\cdots,j_{l},i_{1},\cdots,i_{k},\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}I_{c},I_{d},J_{c},J_{d}$, $K_{c}$ and $K_{d}$ be defined as in Lemma 5.9. Then
@@
Let $a^{(\ell)}$ be the reduced word with the same support as $c^{(\ell)}$, but replacing each $s_{i}a_{i}$ by $a_{i}$; let $b^{(\ell)}$ be defined similarly, and let $w^{\prime}=a^{(\ell)}\circ b^{(\ell)}\circ w$. Note that $w^{\prime}$ consists only of letters of $a$ and $b$. Let $\sigma:I_{c}\rightarrow I_{d}$ be defined as in Lemma 5.9 so that $s_{o}a_{o}=t_{o\sigma}b_{o\sigma}$ and $w_{c,d}$ is the product of letters $s_{o}a_{o}$ where $o\in I_{c}$. We now call upon the fact that all the vertex monoids are left ideal Howson, and we let $X_{o}$ be a finite set of generators of the left ideal $M_{o}a_{o}\cap M_{o}b_{o\sigma}$ for all $o\in I_{c}:=\{o_{1},\cdots,o_{g}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\}$, where $o=\mbox{$\mathbf{s}$}(a_{o})$. We claim that $X$ is a finite set of generators for $\mathscr{I}$, where
$X=\{[x_{o_{1}}\circ\cdots\circ x_{o_{g}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}\circ w^{\prime}]:x_{o_{i}}\in X_{o_{i}},1\leq i\leq g\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\}.$
@@
First, we show that $\mathscr{GP}X\subseteq\mathscr{I}$. Let $x=x_{o_{1}}\circ\cdots\circ x_{o_{g}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}\circ w^{\prime}$ so that $x]\in X$. For $1\leq i\leq g$ choose $p_{o_{i}},q_{o_{i}\sigma}\in M_{o_{i}}$ such that $x_{o_{i}}=p_{o_{i}}a_{o_{i}}=q_{o_{i}\sigma}b_{o_{i}\sigma}$.
@@
We form a word $s_{1}^{\prime}\circ\cdots\circ s_{M}^{\prime}$ with the same support skeleton as $s_{1}\circ\cdots\circ s_{M}$ as follows. If $o\in I_{c}$ then put $s_{o}^{\prime}=p_{o}$; if $o\in J_{c}$ we put $s_{o}^{\prime}=I_{o}$, where $s_{o}\in M_{o}$; if $o\in K_{c}$ let $s_{o}^{\prime}=s_{o}$. Similarly, we form a word $t_{1}^{\prime}\circ\cdots\circ t_{N}^{\prime}$ with the same support skeleton as $t_{1}\circ\cdots\circ t_{N}$ as follows. If $o\sigma\in I_{d}$ then put $t_{o\sigma}^{\prime}=q_{o\sigma}$; if $o\in J_{d}$ we put $t_{o}^{\prime}=I_{o}$, where $t_{o}\in M_{o}$; if $o\in K_{d}$ let $t_{o}^{\prime}=t_{o}$. We now define
$\bar{c}=s_{1}^{\prime}\circ\cdots\circ s_{M}^{\prime}\circ a=s_{1}^{\prime}a_{1}\circ\cdots\circ s_{M}^{\prime}a_{M}\circ a^{M}$
and
$\bar{d}=t_{1}^{\prime}\circ\cdots\circ t_{N}^{\prime}\circ b=t_{1}^{\prime}b_{1}\circ\cdots\circ t_{N}^{\prime}b_{N}\circ b^{N}.$
We call upon Lemma 5.7 to obtain subwords $\bar{c}^{\prime}=\bar{c}_{j_{1}}\circ\cdots\circ\bar{c}_{j_{l}}$ and $\bar{d}^{\prime}=\bar{d}_{i_{1}}\circ\cdots\circ\bar{d}_{i_{k}}$ of $\bar{c}$ and $\bar{d}$ respectively, so that observing that $\bar{c}^{(\ell)}=a^{(\ell)}$ and $\bar{d}^{(\ell)}=b^{(\ell)}$ we have
$\bar{d}^{\prime}\circ\bar{c}\equiv a^{(\ell)}\circ b^{(\ell)}\circ w_{\bar{c},\bar{d}}\circ w\equiv\bar{c}^{\prime}\circ\bar{d}\equiv w_{\bar{c},\bar{d}}\circ w^{\prime}=[x]$
where $w_{\bar{c},\bar{d}}=x_{o_{1}}\circ\cdots\circ x_{o_{g}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}$. This gives that
$x]=[\bar{d}^{\prime}\circ s_{1}^{\prime}\circ\cdots\circ s_{M}^{\prime}\circ a]=[\bar{c}^{\prime}\circ t_{1}^{\prime}\circ\cdots\circ t_{N}^{\prime}\circ b]\in\mathscr{I}$
and, consequently, $\mathscr{GP}X\subseteq\mathscr{I}$.
@@
Now we show that $\mathscr{GP}X\supseteq\mathscr{I}$. To this end, let us return to an arbitrary $z]\in\mathscr{I}$. Then $z]=[s\circ a]=[t\circ b]$ where $s$ and $t$ are obtained as at the start of the proof. Note that it was the existence of any such $[z]$ that determined the support skeleton of $d^{\prime}\circ c$ and $c^{\prime}\circ d$; the support skeleton does not depend upon the choice of $z$ and by Remark 5.8 is unique. For $o_{i}\in I_{c}$ we have that $s_{o_{i}}a_{o_{i}}=t_{o_{i}\sigma}b_{o_{i}\sigma}$ so that $s_{o_{i}}a_{o_{i}}=k_{o_{i}}x_{o_{i}}$ for some $x_{o_{i}}\in X_{o_{i}}$. Put $x_{o_{i}}=p_{o_{i}}a_{o_{i}}$ as above for $1\leq i\leq g\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$ and let $x=x_{o_{1}}\circ\cdots\circ x_{o_{g}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}\circ w^{\prime}$ so that $x]\in X$. Now list the elements of $J_{c}$ and $J_{d}$ respectively as
$J_{c}=\{e_{1},\cdots,e_{h\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}\}\mbox{ and }J_{d}=\{f_{1},\cdots,f_{r}\}.$
@@
$s\circ a]=[u\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\circ(k_{o_{1}}\circ\cdots\circ k_{o_{g}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}\circ s_{e_{1}}\circ\cdots\circ s_{e_{h\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}}}\circ t_{f_{1}}\circ\cdots\circ t_{f_{r}})][x]\in\mathscr{GP}X.$
@@ 4. $\bullet$
We call the pair of reduction functions $(\theta,\psi)$ a parallel pair and say that $([s],[t])\in\mathbf{l}([a])$ belongs to a $(\theta,\psi)$ reduction. Note that there are only finitely many types of parallel pairs in $\mathbf{l}([a])$, since there are only finitely many reduction functions for $a$.
@@
For a parallel pair of reduction functions $(\theta,\psi)$ and $B^{\prime}$ and $A^{\prime}$ defined as above, we let $B^{\prime}_{r}$ be $B^{\prime}$ where we replace any $t_{u_{k}\psi}a_{u_{k}\psi}$ in $B_{1}$ by $a_{u_{k}\psi}$ (so that $B_{r}^{\prime}$ is a product of letters of $a$) and let $A^{\prime}_{r}$ be $A^{\prime}$ where we replace any $s_{i_{l}\theta}a_{i_{l}\theta}$ in $B_{1}$ by $a_{i_{l}\theta}$ (so that $A_{r}^{\prime}$ is a product of letters of $a$). We then set
@@
We have that $([z^{\theta,\psi}_{\theta}],[z^{\theta,\psi}_{\psi}])\in\mathbf{l}([a])$.
@@ 8. Weak left coherency
9. Acknowledgements
We are grateful to the referee for a careful reading of the manuscript and to Levent Dasar for pointing out an error in the original formulation of Proposition 4.8. The first author would like to thank the Department of Mathematics of the University of York for hosting her whilst this work was in development.